OpenStudy (anonymous):

13:52 ration in the simplest form

OpenStudy (anonymous):

ratio

OpenStudy (anonymous):

1:4

OpenStudy (aroub):

13:52 is like saying 13 divided by 52 13/52 =1/4 = 1:4